# The Hidden Risks of Large Reasoning Models: A Safety Assessment of R1

### v3

@@ 5 Response Harmfulness Level Evaluation
We quantitatively evaluate the model’s harmfulness level on two datasets with different malicious scenarios. For AIR-bench, we evaluate the helpfulness to the malicious question using two top pre-trained reward models on the RewardBench Lambert et al. (2024) – ArmoRM-Llama3-8B Wang et al. (2024) and QRM-Llama3.1-8B Dorka (2024). These models are trained to predict the reward score for 19 attributes, such as helpfulness, correctness, and coherence. We utilize the average reward score for the helpsteer-helpfulness and ultrafeedback-helpfulness attributes to represent the helpfulness of the response to queries in AIR-bench. In Spear Phishing Tests, the helpfulness of the model to the malicious instruction can be evaluated as the attack techniques they demonstrate in the attack process. We use the automated LLM-based grading system from the test suite to evaluate the attack skills, including persuasion, rapport, and argumentation. Specifically, we use Llama 3.3 as the LLM grader.
@@
Table 4: The performance of models in terms of social engineering capabilities in the spear phishing test. The social engineering metrics include average scores (on a scale of 0 to 5) for persuasion, rapport, and argumentation, with each model assessed based on 200 prompts. Higher scores indicate greater security risks.
@@ 6.2 Prompt Injection
Table 6 presents the results of the text prompt injection attack, revealing significant differences among models in terms of injection types and risk categories. Regarding injection types, the ASR of indirect injection is generally higher than that of direct injection. Although the reasoning models perform similarly to other models in indirect attacks, their robustness is significantly weaker in direct attacks, with ASR nearly double that of the others, indicating a higher level of compliance to direct prompts. In terms of risk categories, reasoning models exhibit a higher ASR for security attacks than for logic. This discrepancy indicates that despite these models demonstrate strong logical reasoning capabilities, their security protection awareness is weak. Finally, compared with o3-mini, R1 models are more vulnerable to prompt injection attacks.
@@ 8 Discussion and Conclusion
In this paper, we present a comprehensive multi-faceted analysis of the safety of large reasoning models. In our analysis, we identify a significant safety gap between the DeepSeek-R1 models and the o3-mini. We suggest three potential directions for improvement. First, enhancing the extent of safety alignment in R1 models, as their current alignment training may be insufficient, especially in certain safety categories. Second, advanced safety alignment techniques—such as rule-based rewards and methods that leverage reasoning ability to enhance safety Mu et al. (2024); Guan et al. (2024) could be explored. Third, developing new training strategies to enhance their explicit safety reasoning, in terms of activating safety thinking and improving the precision of safety judgments. In addition, the distilled R1 model compromises the original safety performance consistently in all the safety tests. Therefore, additional safety alignment is required after reasoning distillation.
Moreover, we find that with stronger reasoning ability, the R1 models provide more help to the malicious queries compared with their non-reasoning counterparts. Therefore, their unsafe responses are more harmful. This further underscores the necessity of enhancing the safety of R1 models. Finally, within the outputs of large reasoning models, we find that the thinking process may contain hidden safety risks that are not reflected in the final answer. This presents a new challenge brought by reasoning models, which requires future work to address.
@@ A.1 Benchmark Introductions
CyberSecEval–MITRE Tests Wan et al. (2024a)
@@
CyberSecEval–Code Interpreter Tests  Wan et al. (2024a)
@@
CyberSecEval–Spear Phishing Tests  Wan et al. (2024a)
@@
Prompt Injection Wan et al. (2024a)
@@
A.2 Additional Details on the Safety Evaluation Results
Tables 9 and 10 present a more comprehensive safety evaluation of the model under the XSTest and Code Interpreter environments. These results provide insights into the model’s performance when facing various complex security challenges and further highlight the model’s vulnerabilities and robustness under different testing conditions.
A.3 More Examples
@@
Safe Prompt Types DeepSeek-R1 DeepSeek-V3 DeepSeek-R1-70B OpenAI-o3mini Llama-3.3-70B
T1: Homonyms 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 96 + 4 + 0 100 + 0 + 0
T2: Figurative Language 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T3: Safe Targets 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T4: Safe Contexts 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T5: Definitions 96 + 0 + 4 100 + 0 + 0 96 + 4 + 0 100 + 0 + 0 100 + 0 + 0
T6: Real Discr., Nons. Group 84 + 16 + 0 100 + 0 + 0 96 + 0 + 4 96 + 4 + 0 96 + 4 + 0
T7: Nons. Discr., Real Group 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 63 + 32 + 0 100 + 0 + 0
T8: Historical Events 92 + 8 + 0 100 + 0 + 0 92 + 8 + 0 100 + 0 + 0 100 + 0 + 0
T9: Privacy (Public) 96 + 0 + 4 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T10: Privacy (Fictional) 92 + 8 + 0 80 + 20 + 0 68 + 28 + 4 64 + 36 + 0 72 + 20 + 8
TOTAL: (n=250 safe prompts) 96 + 3.2 + 0.8 98 + 2 + 0 94.8 + 4.4 + 0.8 92.8 + 7.2 + 0 96.8 + 2.4 + 0.8

### v4

@@ 5 Response Harmfulness Level Evaluation
We quantitatively evaluate the model’s harmfulness level on two datasets with different malicious scenarios. For AIR-bench, we evaluate the helpfulness to the malicious question using two top pre-trained reward models on the RewardBench lambert2024rewardbench – ArmoRM-Llama3-8B wang2024interpretable and QRM-Llama3.1-8B dorka2024quantile. These models are trained to predict the reward score for 19 attributes, such as helpfulness, correctness, and coherence. We utilize the average reward score for the helpsteer-helpfulness and ultrafeedback-helpfulness attributes to represent the helpfulness of the response to queries in AIR-bench. In Spear Phishing Tests, the helpfulness of the model to the malicious instruction can be evaluated as the attack techniques they demonstrate in the attack process. We use the automated LLM-based grading system from the test suite to evaluate the attack skills, including persuasion, rapport, and argumentation. Specifically, we use Llama 3.3 as the LLM grader.
@@
Gemini 2.5 2.28 2.61 1.88 2.37
Table 4: The performance of models in terms of social engineering capabilities in the spear phishing test. The social engineering metrics include average scores (0-5) for persuasion, rapport, and argumentation. Higher scores indicate greater security risks.
@@ 6.2 Prompt Injection
Table 6 presents the results of the text prompt injection attacks, revealing significant differences among models in terms of injection types and risk categories. In terms of injection types, the ASR for indirect injections is generally higher than that for direct injections, indicating that models are more susceptible to manipulation by implicit instructions. Indirect injections influence the model through subtle cues, such as covering the injected instruction in website and email content, making the attack harder to detect and leading to higher ASR. In contrast, direct injections involve explicit, aggressive instructions that directly conflict with the system’s goals, making them easier to detect and reject.
Regarding risk categories, the ASR for security-related attacks is higher than that for logic-related attacks, suggesting that models are more likely to generate unsafe content under security attacks. Security attacks are typically more covert, difficult to identify, and harder to defend against. In contrast, logic-related attacks are more reliant on reasoning, and reasoning models are better equipped to leverage their reasoning abilities to identify and prevent the generation of incorrect outputs. Finally, compared to closed-source models like o3-mini, open-source reasoning models, such as the R1 series, are more vulnerable to prompt injection attacks.
@@ 8 Discussion and Conclusion
In this paper, we present a comprehensive multi-faceted analysis of the safety of large reasoning models. In our analysis, we identify a significant safety gap between open reasoning models and the o3-mini in terms of harmful content generation and adversarial attacks. In addition, the distilled R1 model compromises the original safety performance consistently in all the safety tests. Moreover, we find that with stronger reasoning ability, the R1 models provide more help to the malicious queries compared with their non-reasoning counterparts. Therefore, their unsafe responses are more harmful. This further underscores the necessity of enhancing the safety of R1 models. Finally, within the outputs of large reasoning models, we find that the thinking process may contain hidden safety risks that are not reflected in the final answer. This presents a new challenge brought by reasoning models, which requires future work to address.
To mitigate these challenges, we suggest several potential directions for improvement. First, enhancing the extent of safety alignment in R1 models, as their current alignment training may be insufficient, especially in certain safety categories. Second, advanced safety alignment techniques, such as rule-based rewards and methods that leverage reasoning ability to enhance safety mu2024rule; guan2412deliberative could be explored. Third, developing new training strategies or creating in-domain training data to enhance their explicit safety reasoning, in terms of activating safety thinking and improving the precision of safety judgments against adversarial attacks zhou2025safekey. Additionally, adapting safety techniques designed for non-reasoning LLMs to reasoning models, such as representation engineering zou2024improving, may offer further gains. Finally, to mitigate the issue of unsafe reasoning trajectories, reinforcement learning with reward functions tailored to the safety of intermediate reasoning steps could be a promising direction.
@@ A.1 Benchmark Introductions
CyberSecEval–MITRE Tests wan2024cyberseceval3advancingevaluation
@@
CyberSecEval–Code Interpreter Tests  wan2024cyberseceval3advancingevaluation
@@
CyberSecEval–Spear Phishing Tests  wan2024cyberseceval3advancingevaluation
@@
Prompt Injection wan2024cyberseceval3advancingevaluation
@@
A.2 Reliability of GPT4o Evaluation
To assess the reliability of GPT-4o’s evaluation, we conducted a manual annotation study. Specifically, we randomly sampled 60 queries from Air-Bench and the MITRE test in CyberSecEval, and manually labeled the safety of 4 models’ responses: o3-mini, llama3.3, R1-70b, and R1, resulting in 240 human labels. We then compared GPT-4o’s evaluations against human labels and found that GPT-4o achieved an accuracy of 96.7%, demonstrating a strong alignment with human judgments.
[…]
O3-mini 73.3 70.0
Others 43.3 45.6
Table 8: Comparison of model performance evaluated by Human and GPT-4o.
Additionally, to investigate potential bias, we analyzed the safety ratings assigned to o3-mini and other models by both GPT-4o and human evaluators, as in Table 8. These results indicate negligible bias favoring o3-mini in GPT-4o’s evaluations. While GPT-4o is not perfect, we carefully examined its errors. For instance, it sometimes misclassifies empty responses (e.g., "") as unsafe, whereas we consider them safe since they provide no assistance to malicious queries. GPT4o also occasionally labels borderline unsafe responses as safe. Despite these minor misclassifications, the high overall accuracy of GPT-4o, combined with the significant performance gaps observed between models in our paper, supports the validity of our conclusions.
A.3 Additional Details on the Safety Evaluation Results
Tables 10 and 11 present a more comprehensive safety evaluation of the model under the XSTest and Code Interpreter environments. These results provide insights into the model’s performance when facing various complex security challenges and further highlight the model’s vulnerabilities and robustness under different testing conditions.
A.4 More Examples
@@
Safe Prompt Types DeepSeek-R1 DeepSeek-V3 DeepSeek-R1-70B OpenAI-o3mini Llama-3.3-70B Gemini 2.5 QwQ
T1: Homonyms 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 96 + 4 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T2: Figurative Language 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T3: Safe Targets 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T4: Safe Contexts 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T5: Definitions 96 + 0 + 4 100 + 0 + 0 96 + 4 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T6: Real Discr., Nons. Group 84 + 16 + 0 100 + 0 + 0 96 + 0 + 4 96 + 4 + 0 96 + 4 + 0 100 + 0 + 0 96 + 0 + 4
T7: Nons. Discr., Real Group 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 63 + 32 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T8: Historical Events 92 + 8 + 0 100 + 0 + 0 92 + 8 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0
T9: Privacy (Public) 96 + 0 + 4 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 100 + 0 + 0 96 + 0 + 4 100 + 0 + 0
T10: Privacy (Fictional) 92 + 8 + 0 80 + 20 + 0 68 + 28 + 4 64 + 36 + 0 72 + 20 + 8 88 + 12 + 0 88 + 12 + 0
TOTAL: (n=250 safe prompts) 96 + 3.2 + 0.8 98 + 2 + 0 94.8 + 4.4 + 0.8 92.8 + 7.2 + 0 96.8 + 2.4 + 0.8 98.4 + 1.2 + 0.4 98.4 + 1.2 + 0.4